### v1

# Constraints on inflationary decoherence from attractor model

Abstract
The quantum-to-classical transition of inflationary perturbations remains a fundamental challenge in cosmology, and quantum decoherence is one of the proposed solutions to this issue. By treating primordial quantum perturbations as an open quantum system interacting with environmental degrees of freedom during inflation, decoherence can be described by the Lindblad master equation. Lindblad equation modifies the dynamical evolution of the primordial quantum perturbations and affects the power spectrum of the curvature perturbations, resulting in changes in the primordial power spectrum. In this paper, we examine the decoherence process of a polynomial attractor model that includes an ultra-slow-roll plateau. We compare our findings with previous results from slow-roll analyses. We numerically compute the corrections to the power spectrum due to decoherence. The results show significant modifications only on large scales, with a peak appearing at the minimum of the power spectrum. Using CMB constraints on the spectral index and tensor-to-scalar ratio, and requiring complete decoherence for relevant scales by the end of inflation, we constrain the interaction parameter: $10^{-28}\text{Mpc}^{-1}<k_{\gamma}<0.042\text{Mpc}^{-1}$.
I Introduction
The inflationary theory is one of the most successful theories in cosmology today. It assumes that the universe experienced a rapid expansion in its very early stages, which naturally addresses the horizon and flatness problems of the universe [1, 2, 3, 4, 5]. The primordial quantum fluctuations during the inflationary epoch can provide the primordial density perturbations that lead to the formation of large-scale structures (LSS) in the universe and anisotropies in the Cosmic Microwave Background (CMB) radiation [6, 7, 8, 9, 10, 11], which is supported by observations [12].
However, inflationary theory faces several unresolved challenges, with one of the most important being the quantum-to-classical transition problem [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. The observational quantities in modern cosmology, including the temperature anisotropies of the CMB and the statistical properties of the LSS, demonstrate classical behavior. The question is how the primordial quantum fluctuations that seeded these structures made the transition to classical ones. A central issue arises from modeling primordial quantum perturbations as isolated systems. In reality, these perturbations interact with environmental degrees of freedom, leading to decoherence [38, 39, 40] that effectively converts quantum fluctuations into classical stochastic density inhomogeneities.
The dynamics of an open quantum system are described by the Lindblad master equation [41], which describes the time evolution of a quantum system interacting with an external environment. Under the environmental interactions, decoherence suppresses the off-diagonal elements of the system’s density matrix, driving it to a diagonal state [42, 43, 44]. This equation leads to a correction to the outcomes of the primordial perturbations, thereby affecting their power spectrum. In this paper, we consider the effect of decoherence and numerically compute the modified power spectrum of a inflationary model with an ultra-slow-roll stage. Moreover, we constrain the interaction parameter from the CMB observations.
The process of decoherence takes time and is considered complete when the off-diagonal elements of the quantum system’s density matrix become negligible. The decoherence time depends on the strength of the interaction between the system and its environment: stronger coupling leads to shorter timescales. On cosmological scales, the primordial quantum perturbations undergo complete decoherence by the end of inflation, thereby transitioning to a classical state. In this paper, we require that decoherence occurs sufficiently at the scale of interest, which imposes a lower bound constraint on the interaction parameter.
This paper is organized as follows. In Sec. II, we present the Lindblad equation that the quantum perturbations satisfy during inflation, along with its exact solution under the assumption of linear interactions. In Sec. III, we provide the parameters of our model and present the results of numerical calculations regarding the impact of decoherence, as well as constraints on the interaction strength. We summarize this paper in Sec. IV.
II Lindblad equation and the modified power spectrum
The inflationary curvature perturbation is described by the gauge-invariant Mukhanov-Sasaki scalar variable $v(\eta,\bm{x})$ [6]. The free Hamiltonian of the perturbation in Fourier space is
$\hat{H}_{\bm{k}}=\frac{1}{2}\int_{\mathbb{R}^{3}}\mathrm{d}^{3}\bm{k}\left[\omega^{2}(\eta,\bm{k})\hat{v}_{\bm{k}}\hat{v}^{\dagger}_{\bm{k}}+\hat{p}_{\bm{k}}\hat{p}^{\dagger}_{\bm{k}}\right],$ (1)
where $\hat{v}_{k}$ is the Mukhanov-sasaki variable in Fourier space and $\hat{p}_{k}$ is its conjugate momentum, which obeys $\hat{p}_{k}=\hat{v}_{k}^{\prime}$. A prime denotes a derivative with respect to conformal time $\eta$. The frequency of mode $\bm{k}$ is given by
$\omega^{2}(\eta,\bm{k})=\bm{k}^{2}-\frac{(\sqrt{\epsilon_{1}}a)^{\prime\prime}}{\sqrt{\epsilon_{1}}a},$ (2)
where $\epsilon_{1}$ is the first slow-roll parameter, $\epsilon_{1}=-\dot{H}/H^{2}$, with $H$ representing the Hubble parameter and a dot indicating differentiation with respect to time. The Hamiltonian (1) describes a isolated system. However, the primordial perturbations could interact with other degrees of freedom in the early universe, making it more realistic to consider the perturbations as an open quantum system [32, 35].
For an open quantum system that interacts with the environment, we can describe the total system in the Hilbert space $\mathcal{E}=\mathcal{E}_{S}\otimes\mathcal{E}_{E}$, where the total Hamiltonian is
$\hat{H}=\hat{H}_{0}+\hat{H}_{\mathrm{int}}=\hat{H}_{S}\otimes\hat{I}_{E}+\hat{I}_{S}\otimes\hat{H}_{E}+g\hat{H}_{\mathrm{int}},$ (3)
where $H_{S}$ is the Hamiltonian of the system, which acts in $\mathcal{E}_{S}$, and $H_{E}$ is the Hamiltonian of the environment, which acts in $\mathcal{E}_{E}$, $H_{\mathrm{int}}$ is the interaction Hamiltonian and $g$ is the coupling constant. Suppose that the system and the environment only couple through local interactions, then the interaction Hamiltonian can be written as
$\hat{H}_{\mathrm{int}}=\int\mathrm{d}^{3}\bm{x}\hat{A}\otimes\hat{B},$ (4)
where $\hat{A}$ acts as a local operator within the system’s subsystem, while $\hat{B}$ represents a local operator associated with the environmental component. The symbol $\bm{x}$ represents the conformal spatial coordinate. According to the reasons mentioned in [32, 35], we can consider $\hat{A}$ to have the form
$\hat{A}=\hat{v}^{n},$ (5)
where $n$ is a constant. By tracing out the environmental degrees of freedom from the total density matrix $\hat{\rho}_{\text{tot}}$, we can derive the reduced density matrix for the system $\hat{\rho}_{S}$
$\hat{\rho}_{S}=\mathrm{Tr}_{E}\,\hat{\rho}_{\text{tot}}.$ (6)
Using the Born and Markov approximations, we can derive the equation of motion of $\hat{\rho}_{S}$, i.e. the Lindblad equation [41, 42, 32]
$\frac{\mathrm{d}\hat{\rho}_{S}}{\mathrm{d}\eta}=i\left[\hat{\rho}_{S},\hat{H}_{S}\right]-\frac{\gamma}{2}\int\mathrm{d}^{3}\bm{x}\,\mathrm{d}^{3}{\bm{y}}\,C_{B}\left({\bm{x}},{\bm{y}}\right)\left[\left[\hat{\rho}_{S},\hat{A}(\bm{y})\right],\hat{A}(\bm{x})\right],$ (7)
where $\gamma=2g^{2}\eta_{c}$ with $\eta_{c}$ being the conformal auto-correlation time of $\hat{B}$, and $C_{B}$ is the same-time correlation function of $\hat{B}$, given by $C_{B}(\bm{x},\bm{y})=\text{Tr}_{E}[\rho_{E}\hat{B}(\eta,\bm{x})\hat{B}(\eta,\bm{y})]$. We assume that the time-dependent parameter $\gamma$ depends on the scale factor as a power law [32]
$\gamma=\gamma_{*}\left(\frac{a}{a_{*}}\right)^{p},$ (8)
where * denotes a reference time when the pivot scale $k_{*}=0.05\mathrm{Mpc}^{-1}$ crosses the Hubble radius, and $p$ is a free parameter. Assuming the environment to be isotropic and statistically homogeneous, $C_{B}(\bm{x},\bm{y})$ can be defined as a top hat function [32]
$C_{B}(\bm{x},\bm{y})=\bar{C}_{B}\Theta\left(\frac{a|\bm{x}-\bm{y}|}{\ell_{\mathrm{E}}}\right),$ (9)
where $\ell_{\mathrm{E}}$ is a characteristic physical correlation length and
$\Theta(x)=\begin{cases}1,&\mathrm{if}\ x<1\\ 0,&\mathrm{otherwise}\end{cases}.$ (10)
Expanding Eq. (9) in Fourier space, we can obtain
$\tilde{C}_{B}(\bm{k})\approx\sqrt{\frac{2}{\pi}}\frac{\bar{C}_{B}\ell^{3}_{\mathrm{E}}}{3a^{3}}\Theta\left(\frac{k\ell_{\mathrm{E}}}{a}\right).$ (11)
Considering the case where the environment consists of a heavy scalar field $\varphi$ with mass $M\gg H$, and the interaction Hamiltonian is
$\hat{H}_{\mathrm{int}}=\lambda\mu^{4-n-m}\int\mathrm{d}^{3}\bm{x}\sqrt{-g}\,\hat{\phi}^{n}(\eta,\bm{x})\hat{\varphi}^{m}(\eta,\bm{x}),$ (12)
where the parameter $\mu$ is defined as a constant mass scale and $\hat{\phi}=\hat{v}/a$. In this case, the parameters can be represented as [32]
$\bar{C}_{B}$ =$=\left\{\left(2m-1\right)!!-\sigma\left(m\right)\left[\left(m-1\right)!!\right]^{2}\right\}\left(\frac{37}{504\pi^{2}}\frac{H^{6}}{M^{4}}\right)^{m}\,,$ (13)
$a\eta_{\mathrm{c}}$ $=\ell_{E}=2\sqrt{2}\sqrt{\frac{\left(2m-1\right)!!-\sigma\left(m\right)\left[\left(m-1\right)!!\right]^{2}}{m^{2}\left(2m-3\right)!!}}\frac{1}{M},$ (14)
$\gamma$ $=4\sqrt{2}\sqrt{\frac{(2m-1)!!-\sigma(m)[(m-1)!!]^{2}}{m^{2}(2m-3)!!}}\frac{\lambda^{2}}{M}\mu^{8-2n-2m}a^{7-2n},$ (15)
where !!double-factorial!!!! denotes the double factorial and
$\sigma(m)=\begin{cases}1,\quad m\,\,\mathrm{is\,\,even}\\ 0,\quad m\,\,\mathrm{is\,\,odd}\end{cases}.$ (16)
In order to solve the Lindblad equation (7), we decompose $\hat{v}_{\bm{k}}$ and $\hat{p}_{\bm{k}}$ into their real and imaginary parts
$\hat{v}_{\bm{k}}=\frac{1}{\sqrt{2}}\left(\hat{v}_{\bm{k}}^{\mathrm{R}}+i\hat{v}_{\bm{k}}^{\mathrm{I}}\right),\quad\hat{p}_{\bm{k}}=\frac{1}{\sqrt{2}}\left(\hat{p}_{\bm{k}}^{\mathrm{R}}+i\hat{p}_{\bm{k}}^{\mathrm{I}}\right).$ (17)
For a real operator $\hat{v}(\eta,\bm{x})$, we have $\hat{v}_{-\bm{k}}=\hat{v}^{\dagger}_{\bm{k}}$, which implies $\hat{v}^{\mathrm{R}}_{-\bm{k}}=\hat{v}^{\mathrm{R}}_{\bm{k}}$ and $\hat{v}^{\mathrm{I}}_{-\bm{k}}=-\hat{v}^{\mathrm{I}}_{\bm{k}}$. These relations show that not all $\hat{v}_{\bm{k}}$ operators are independent degrees of freedom, we can calculate the variables $v_{\bm{k}}^{\mathrm{R}}$ and $v_{\bm{k}}^{\mathrm{I}}$ in $\bm{k}\in\mathbb{R}^{3+}$. Under the condition of linear interaction, the free Hamiltonian (1) can be written as
$\hat{H}_{\bm{k}}^{s}=\frac{1}{2}\int_{\mathbb{R}^{3+}}\mathrm{d}^{3}\bm{k}\left[\omega^{2}(\eta,\bm{k})(\hat{v}_{\bm{k}}^{s})^{2}+(\hat{p}_{\bm{k}}^{s})^{2}\right],$ (18)
[…]
$\hat{\rho}_{S}=\prod_{\bm{k}\in\mathbb{R}^{3+}}\prod_{s=\mathrm{R},\mathrm{I}}\hat{\rho}_{\bm{k}}^{s}.$ (19)
then we can obtain the Lindblad equation in the Fourier space for linear interaction [32]
$\frac{\mathrm{d}\hat{\rho}_{\bm{k}}^{s}}{\mathrm{d}\eta}=-i\left[\hat{\mathcal{H}}_{\bm{k}}^{s},\hat{\rho}_{\bm{k}}^{s}\right]-\frac{\gamma}{2}(2\pi)^{3/2}\tilde{C}_{B}({\bm{k}})\left[\hat{v}_{\bm{k}}^{s},\left[\hat{v}_{\bm{k}}^{s},\hat{\rho}_{\bm{k}}^{s}\right]\right].$ (20)
Note that Eq. (20) is valid only for linear interaction. We can define a combined interaction parameter as follows
$k_{\gamma}\equiv\sqrt{\frac{8\pi}{3}\bar{C}_{B}{}_{*}\ell_{E}^{3}\frac{\gamma_{*}}{a_{*}^{3}}},$ (21)
which combines all the parameters of interaction and possesses the dimension of comoving wavenumber.
Then we define the eigenvectors $|v_{\rm k}^{s}\rangle$ of the operator $\hat{v}_{\rm k}^{s}$, which satisfy equation $\hat{v}_{\rm k}^{s}|v_{\mathbf{k}}^{s}\rangle=v_{\rm k}^{s}|v_{\rm k}^{s}\rangle$. By applying the bra $\langle v_{\bm{k}}^{s,(1)}|$ and the ket $|v_{\bm{k}}^{s,(2)}\rangle$ to the Lindblad equation Eq. (20), the Lindblad equation can be solved exactly, the solution is given by [32]
$\left\langle v_{\bm{k}}^{s,(1)}\right|\hat{\rho}_{\bm{k}}^{s}\left|v_{\bm{k}}^{s,(2)}\right\rangle=\frac{\left(2\pi\right)^{-1/2}}{\sqrt{\left|v_{\bm{k}}\right|^{2}+\mathcal{J}_{\bm{k}}}}\mathrm{exp}\left\{-\frac{{v^{s,(2)}_{\bm{k}}}^{2}+{v^{s,(1)}_{\bm{k}}}^{2}+i{\left|v_{\bm{k}}\right|^{2}}^{\prime}\left[{v^{s,(2)}_{\bm{k}}}^{2}-{v^{s,(1)}_{\bm{k}}}^{2}\right]}{4\left(\left|v_{\bm{k}}\right|^{2}+\mathcal{J}_{\bm{k}}\right)}\right\}$
×$\qquad\qquad\times\exp\Biggl{\{}-\frac{\left[v^{s,(2)}_{\bm{k}}-v^{s,(1)}_{\bm{k}}\right]^{2}}{2\left(\left|v_{\bm{k}}\right|^{2}+\mathcal{J}_{\bm{k}}\right)}\left(\mathcal{I}_{\bm{k}}\mathcal{J}_{\bm{k}}-\mathcal{K}_{\bm{k}}^{2}+\left|v_{\bm{k}}^{\prime}\right|^{2}\mathcal{J}_{\bm{k}}+\left|v_{\bm{k}}\right|^{2}\mathcal{I}_{\bm{k}}-{\left|v_{\bm{k}}\right|^{2}}^{\prime}\mathcal{K}_{\bm{k}}\right)$
$\qquad\qquad-\frac{i\mathcal{K}_{\bm{k}}}{2\left(\left|v_{\bm{k}}\right|^{2}+\mathcal{J}_{\bm{k}}\right)}\left[{v^{s,(2)}_{\bm{k}}}^{2}-{v^{s,(1)}_{\bm{k}}}^{2}\right]\Biggr{\}},$ (22)
where $\mathcal{I}_{\bm{k}}$, $\mathcal{J}_{\bm{k}}$ and $\mathcal{K}_{\bm{k}}$ are defined by
$\mathcal{I}_{\bm{k}}\left(\eta\right)$ $\equiv 4\left(2\pi\right)^{3/2}\int_{-\infty}^{\eta}\mathrm{d}\eta^{\prime}\gamma\left(\eta^{\prime}\right)\tilde{C}_{B}\left(\bm{k},\eta^{\prime}\right)\mathrm{Im}^{2}\left[v_{\bm{k}}\left(\eta^{\prime}\right){v_{\bm{k}}^{*}}^{\prime}\left(\eta\right)\right]\,,$ (23)
$\mathcal{J}_{\bm{k}}\left(\eta\right)$ $\equiv 4\left(2\pi\right)^{3/2}\int_{-\infty}^{\eta}\mathrm{d}\eta^{\prime}\gamma\left(\eta^{\prime}\right)\tilde{C}_{B}\left(\bm{k},\eta^{\prime}\right)\mathrm{Im}^{2}\left[v_{\bm{k}}\left(\eta^{\prime}\right){v_{\bm{k}}^{*}}\left(\eta\right)\right]\,,$ (24)
$\mathcal{K}_{\bm{k}}\left(\eta\right)$ $\equiv 4\left(2\pi\right)^{3/2}\int_{-\infty}^{\eta}\mathrm{d}\eta^{\prime}\gamma\left(\eta^{\prime}\right)\tilde{C}_{B}\left(\bm{k},\eta^{\prime}\right)\mathrm{Im}\left[v_{\bm{k}}\left(\eta^{\prime}\right){v_{\bm{k}}^{*}}^{\prime}\left(\eta\right)\right]\mathrm{Im}\left[v_{\bm{k}}\left(\eta^{\prime}\right){v_{\bm{k}}^{*}}\left(\eta\right)\right]\,,$ (25)
where $v_{\bm{k}}(\eta)$ is the solution of the Mukhanov-Sasaki equation. Using Eq. (II) and the relation $\langle\hat{O}\rangle=\mathrm{Tr}(\hat{\rho}_{v}\hat{O})$ with $\hat{O}=\hat{v}_{\bm{k}}^{2}$, we can obtain the two-point correlation function
$P_{vv}(k)=\left|v_{\bm{k}}\right|^{2}+\mathcal{J}_{\bm{k}},$ (26)
where $|v_{\bm{k}}|^{2}$ is the standard two-point correlation and $\mathcal{J}_{\bm{k}}$ is the correction from the interaction with the environment. The modified power spectrum is
$\mathcal{P}_{\zeta}=\frac{k^{3}}{2\pi^{2}}\frac{P_{vv}}{2a^{2}\epsilon_{1}}=\mathcal{P}_{\zeta}|_{\mathrm{standard}}(1+\Delta\mathcal{P}_{\bm{k}}),$ (27)
where we take $M_{\mathrm{Pl}}^{2}=1$. On the other hand, for linear interaction, the tensor power spectrum remains unchanged by the corrections in the Lindblad equation [35]. Therefore, we can utilize the spectral index and the tensor-to-scalar ratio to constrain $k_{\gamma}$ for linear interaction.
The interaction with the environment tends to suppress the off-diagonal elements of the density matrix when expressed in the basis of the eigenstates of the interaction operator [42, 32]. From the exact solution of the density matrix Eq. (II), we can study decoherence during inflation. The decoherence parameter $\delta_{\bm{k}}$ defined in [32] represents the reduction in the off-diagonal elements of the density matrix due to environmental interactions, and is defined as follows
$\delta_{\bm{k}}(\eta)\equiv\mathcal{I}_{\bm{k}}\mathcal{J}_{\bm{k}}-\mathcal{K}_{\bm{k}}^{2}+\left|v_{\bm{k}}^{\prime}\right|^{2}\mathcal{J}_{\bm{k}}+\left|v_{\bm{k}}\right|^{2}\mathcal{I}_{\bm{k}}-{\left|v_{\bm{k}}\right|^{2}}^{\prime}\mathcal{K}_{\bm{k}}\,.$ (28)
It is obvious that $\delta_{\bm{k}}$ is related to $k_{\gamma}$. When $\delta_{\bm{k}}\gg 1$, decoherence is expected to be complete. We determine that by the end of inflation, the decoherence of the scales of interest should be finished, which in turn constrains $k_{\gamma}$.
In the study referenced as [32], the researchers took into account the constant $\epsilon_{1}$ and the constant Hubble parameter $H$. In this paper, we will compute the results for a concrete model with an ultra-slow-roll stage. We compare the results of our model’s numerical computations with the analytical results reported in [32]. Then, we constraint $k_{\gamma}$ using the polynomial attractor model.
III Models and numerical results
We choose the modified polynomial attractor model in [45]
$V(\psi)=V_{0}[1+c_{4}f^{-1}(\psi)+c_{1}f(\psi)+c_{2}f^{2}(\psi)+c_{3}f^{3}(\psi)]^{2},$ (29)
where $f(\psi)=1/\psi^{3}$, the parameters of the model are given by Table. 1.
From Eq. (8) and (15), we can see that for linear interaction,
$\gamma(\eta^{\prime})=\gamma_{*}\left(\frac{a(\eta^{\prime})}{a_{*}}\right)^{5}.$ (30)
From Eq. (11) and (13), we have
$\tilde{C}_{B}(\bm{k},\eta^{\prime})=\left(\frac{H(\eta^{\prime})}{H_{*}}\right)^{6m}\left(\frac{a_{*}^{3}}{a(\eta^{\prime})^{3}}\right)\sqrt{\frac{2}{\pi}}\frac{\bar{C}_{B}^{*}\ell_{E}^{3}}{3a_{*}^{3}}\Theta\left(\frac{k\ell_{E}}{a(\eta^{\prime})}\right),$ (31)
where
$\bar{C}_{B}^{*}=\left\{\left(2m-1\right)!!-\sigma\left(m\right)\left[\left(m-1\right)!!\right]^{2}\right\}\left(\frac{37}{504\pi^{2}}\frac{H_{*}^{6}}{M^{4}}\right)^{m}\,.$ (32)
Combine Eq. (24), (30)and (31), we have
$4(2\pi)^{3/2}\gamma(\eta^{\prime})\tilde{C}_{B}(\bm{k},\eta^{\prime})$ $=2\left(\frac{a(\eta^{\prime})}{a_{*}}\right)^{2}\left(\frac{H(\eta^{\prime})}{H_{*}}\right)^{6m}\frac{8\pi}{3}\frac{\bar{C}_{B}^{*}\ell_{E}^{3}\gamma_{*}}{a_{*}^{3}}\Theta\left(\frac{k\ell_{E}}{a(\eta^{\prime})}\right)$
$=2\left(\frac{a(\eta^{\prime})}{a_{*}}\right)^{2}\left(\frac{H(\eta^{\prime})}{H_{*}}\right)^{6m}k_{\gamma}^{2}\,\Theta\left(\frac{k\ell_{E}}{a(\eta^{\prime})}\right).$ (33)
The remaining sector of Eq. (24) can be written as
$\mathrm{Im}^{2}\left[v_{\bm{k}}(\eta^{\prime})v_{\bm{k}}^{*}(\eta)\right]=$$\left\{\mathrm{Im}\left[\left(v^{\mathrm{R}}_{\bm{k}}(\eta^{\prime})+iv^{\mathrm{I}}_{\bm{k}}(\eta^{\prime})\right)\left(v^{\mathrm{R}}_{\bm{k}}(\eta)-iv^{\mathrm{I}}_{\bm{k}}(\eta)\right)\right]\right\}^{2}$
$=$ $\left(v^{\mathrm{R}}_{\bm{k}}(\eta^{\prime})\right)^{2}\left(v^{\mathrm{I}}_{\bm{k}}(\eta)\right)^{2}+\left(v^{\mathrm{I}}_{\bm{k}}(\eta^{\prime})\right)^{2}\left(v^{\mathrm{R}}_{\bm{k}}(\eta)\right)^{2}$
$-2v^{\mathrm{R}}_{\bm{k}}(\eta^{\prime})v^{\mathrm{I}}_{\bm{k}}(\eta^{\prime})v^{\mathrm{R}}_{\bm{k}}(\eta)v^{\mathrm{I}}_{\bm{k}}(\eta),$ (34)
We take $m=3$ so that $\mu$ in Eq. (15) vanishes. In order to compare with [32], we take $k_{\gamma}=5\times 10^{-5}\text{Mpc}^{-1}$ and $H_{*}\ell_{E}=10^{-3}$. The top hat function in Eq. (III) can be written as
$\Theta\left(\frac{k\ell_{E}}{a(\eta^{\prime})}\right)=\Theta\left(\frac{k}{a(\eta^{\prime})H(\eta^{\prime})}\frac{H(\eta^{\prime})}{H_{*}}H_{*}\ell_{E}\right).$ (35)
Then we can calculate the modified power spectrum numerically.
Models $V_{0}$ $\delta$ $c_{1}$ $c_{4}$ $\psi_{*}$ $N$ $n_{s}$ $r$ $k_{\text{peak}}$
A1 8.33×$\times 10^{-10}$ 1.1007×$\times 10^{-3}$1.07 3.2×$\times 10^{-4}$ 3.75 63.2 0.9658 0.025 $3.0\times 10^{8}$
Table 1: Parameters of the model A1.
Figure 1: The black dashed line is the modified power spectrum, the red dotted line is the slow-roll result of $\Delta P_{k}$, the blue solid line is the numerical result of $\Delta P_{k}$ calculated from model A1.
We calculate the modified term $\Delta P_{k}$ of the attractor model A1. The results of the modified power spectrum and $\Delta P_{k}$ are shown in Fig. 1. From Fig. 1, we see that at the pivot scale $k_{*}$, both the slow-roll model and ultra-slow-roll model have the same value of $\Delta P_{k}$. However, as the Hubble parameter gradually decreases, $\Delta P_{k}$ also decreases. We also find that at the minimum of the power spectrum, $\Delta P_{k}$ has a peak in the order of $10^{-6}$, but the correction near the peak is negligible ($\sim 10^{-8}$). This means that the correction due to decoherence almost does not change the abundance of primordial black holes and the scalar induced gravitational waves.
Figure 2: $n_{s}$ and $r$ with respect to $k_{\gamma}$. The red solid lines is the numerical results from model A1, the black dashed lines is the analytical results for slow-roll.
Figure 3: The spectral index and tensor-to-scalar ratio vary with $k_{\gamma}$. The red solid line is the numerical result, the black dashed line is the analytical result, the black dash dot line is the restriction of $n_{s}$ and $r$ at $1\sigma$ and $2\sigma$ from Planck18. As $k_{\gamma}$ increases, the spectral index and tensor-to-scalar ratio decrease.
Figure 4: Decoherence parameter $\delta_{\bm{k}}$ as a function of e-folds $N$ at $k=k_{*}$. The black dashed line is the result of slow-roll model, the red dotted line is the result of model A1. The vertical solid line indicates the e-folds value at the starting point of $\delta_{k}$ evolution.
From Fig. 1, we find that the correction at pivot scale is more significant, which means that we can restrict $k_{\gamma}$ using the spectral tilt $n_{s}$ and the tensor-to-scalar ratio $r$ at pivot scale. The analytical results for $n_{s}$ and $r$ in the case of linear interaction are [32]
$n_{s}=n_{s}|_{\mathrm{standard}}-\frac{\alpha}{1+\alpha}(6m-2)\epsilon_{1*},$ (36)
and
$r=\frac{r|_{\mathrm{standard}}}{1+\alpha},\quad\alpha=\frac{\pi}{6}\frac{k_{\gamma}^{2}}{k_{*}^{2}}.$ (37)
We perform numerical computations to determine the relationships between $k_{\gamma}$, $n_{s}$ and $r$. The results are shown in Fig. 2. We can see that when $k_{\gamma}>10^{-2}\,\text{Mpc}^{-1}$, the results begin to change significantly, when $k_{\gamma}>0.2\text{Mpc}^{-1}$, the changes in the results are minimal. The tensor-to-scalar ratios are well-aligned, whereas the spectral tilts deviate at large $k_{\gamma}$. We plot the results in the $n_{s}-r$ plane to restrict the upper bound of $k_{\gamma}$, the results are shown in Fig. 3. Starting from the initial point, as $k_{\gamma}$ grows, both $n_{s}$ and $r$ diminish. When $k_{\gamma}=0.026\,\text{Mpc}^{-1}$ and $k_{\gamma}=0.042\,\text{Mpc}^{-1}$, the results are within the $1\sigma$ and $2\sigma$ confidence intervals of the observation, respectively. This is the upper bound of $k_{\gamma}$ given by model A1.
We then calculate the decoherence parameter $\delta_{\bm{k}}$ at pivot scale and compare with the result in [32]. The results are shown in Fig. 4. Since model A1 is in the slow-roll stage at the pivot scale, the analytical and numerical results are almost the same. After $N-N_{*}=4$, $\delta_{\bm{k}}$ increase exponentially. We require that for the scales $k_{*}$ and $k_{\mathrm{peak}}$, decoherence is complete. For model A1, this imposes a constraint on $k_{\gamma}$ such that $k_{\gamma}$ must be much greater than $10^{-28}\,\text{Mpc}^{-1}$. This is a rather relaxed constraint.
IV Conclusion
In this paper, we examine the effects of decoherence on the power spectrum of a specific inflationary model that incorporates an ultra-slow-roll phase. Decoherence arises from the interaction of a quantum system with its environment and is described by the Lindblad equation. Based on the solutions to the Lindblad equation for linear interactions provided by J. Martin et al. [32], we numerically computed the corrections to the power spectrum of a polynomial attractor model due to quantum decoherence and constrained the interaction parameters $k_{\gamma}$ using observations. We found that decoherence significantly alters the power spectrum only on large scales, and it is consistent with previous work at the pivot scales. As the scale decreases, the corrections to the power spectrum due to decoherence also diminish, and at the trough of the power spectrum, the corrections will result in a peak. We also calculated the relationship between the spectral index and the tensor-to-scalar ratio at the CMB with $k_{\gamma}$, and using the observational data from Planck 2018, we determined an upper limit for $k_{\gamma}$ of $0.042$  Mpc-1. We calculate the decoherence parameter $\delta_{k}$, which is used to characterize the degree of completion of decoherence; decoherence is considered to be complete when $\delta_{k}\gg 1$. We computed the evolution of $\delta_{k}$ over time at the pivot scale, and the results are roughly consistent with the analysis of the slow-roll regime. We require that decoherence be completed at the peak scale of the power spectrum by the end of inflation, which will provide a lower limit for the interaction parameter $k_{\gamma}$. Our model suggests a lower limit of $10^{-28}\,\text{Mpc}^{-1}$. Since exact solutions are only provided for linear interactions, the aforementioned numerical calculations must satisfy the conditions of linear interaction.

### v2

@@
-Constraints on inflationary decoherence from attractor model
+Observational constraints on inflationary decoherence with polynomial attractor model
 Abstract
-The quantum-to-classical transition of inflationary perturbations remains a fundamental challenge in cosmology, and quantum decoherence is one of the proposed solutions to this issue.
-By treating primordial quantum perturbations as an open quantum system interacting with environmental degrees of freedom during inflation, decoherence can be described by the Lindblad master equation.
-Lindblad equation modifies the dynamical evolution of the primordial quantum perturbations and affects the power spectrum of the curvature perturbations, resulting in changes in the primordial power spectrum.
-In this paper, we examine the decoherence process of a polynomial attractor model that includes an ultra-slow-roll plateau.
-We compare our findings with previous results from slow-roll analyses.
-We numerically compute the corrections to the power spectrum due to decoherence.
-The results show significant modifications only on large scales, with a peak appearing at the minimum of the power spectrum.
-Using CMB constraints on the spectral index and tensor-to-scalar ratio, and requiring complete decoherence for relevant scales by the end of inflation, we constrain the interaction parameter: $10^{-28}\text{Mpc}^{-1}<k_{\gamma}<0.042\text{Mpc}^{-1}$.
+The quantum-to-classical transition of inflationary perturbations remains an unresolved fundamental problem, and quantum decoherence is one of the promising solutions.
+By considering quantum perturbations during inflation as an open quantum system interacting with its environment, quantum decoherence can be described by the Lindblad equation.
+This formalism modifies the evolution of primordial quantum perturbations and consequently alters the power spectrum of curvature perturbations, leading to observable consequences.
+In this paper, we examine the decoherence process of a polynomial attractor model featuring an ultra-slow-roll stage, extending previous analyses limited to slow-roll scenarios.
+We numerically compute the correction to the power spectrum due to quantum decoherence, and the results show significant modification only on large scales, with a peak generated by the decoherence correction at the minimum of the power spectrum.
+Using observational constraints on the scalar spectral index and the tensor-to-scalar ratio, and requiring complete decoherence for relevant scales by the end of inflation, we obtain the constraint on the interaction parameter as $10^{-17}\text{Mpc}^{-1}<k_{\gamma}<0.061\text{Mpc}^{-1}$.
 I Introduction
-The inflationary theory is one of the most successful theories in cosmology today.
-It assumes that the universe experienced a rapid expansion in its very early stages, which naturally addresses the horizon and flatness problems of the universe [1, 2, 3, 4, 5].
-The primordial quantum fluctuations during the inflationary epoch can provide the primordial density perturbations that lead to the formation of large-scale structures (LSS) in the universe and anisotropies in the Cosmic Microwave Background (CMB) radiation [6, 7, 8, 9, 10, 11], which is supported by observations [12].
-However, inflationary theory faces several unresolved challenges, with one of the most important being the quantum-to-classical transition problem [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37].
-The observational quantities in modern cosmology, including the temperature anisotropies of the CMB and the statistical properties of the LSS, demonstrate classical behavior.
-The question is how the primordial quantum fluctuations that seeded these structures made the transition to classical ones.
-A central issue arises from modeling primordial quantum perturbations as isolated systems.
+The inflationary scenario is one of the most successful progresses in cosmology.
+It assumes that the universe experienced a rapid expansion in its very early stages, which naturally addresses the monopole, horizon and flatness problems [1, 2, 3, 4, 5].
+Quantum fluctuations during inflation generate primordial density perturbations that seed the formation of large-scale structures (LSS) observed today and imprint temperature anisotropies in the cosmic microwave background (CMB) radiation [6, 7, 8, 9, 10, 11].
+However, the inflationary scenario faces several unresolved challenges, with a significant one being the quantum-to-classical transition problem [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37].
+The observables such as the temperature anisotropies of the CMB and the statistical properties of the LSS, are inherently classical in nature.
+This raises the critical question of how the primordial quantum fluctuations that seeded these structures became classical.
+The core issue stems from conventional treatments that model primordial perturbations as isolated quantum systems.
 In reality, these perturbations interact with environmental degrees of freedom, leading to decoherence [38, 39, 40] that effectively converts quantum fluctuations into classical stochastic density inhomogeneities.
-The dynamics of an open quantum system are described by the Lindblad master equation [41], which describes the time evolution of a quantum system interacting with an external environment.
-Under the environmental interactions, decoherence suppresses the off-diagonal elements of the system’s density matrix, driving it to a diagonal state [42, 43, 44].
-This equation leads to a correction to the outcomes of the primordial perturbations, thereby affecting their power spectrum.
-In this paper, we consider the effect of decoherence and numerically compute the modified power spectrum of a inflationary model with an ultra-slow-roll stage.
-Moreover, we constrain the interaction parameter from the CMB observations.
-The process of decoherence takes time and is considered complete when the off-diagonal elements of the quantum system’s density matrix become negligible.
-The decoherence time depends on the strength of the interaction between the system and its environment: stronger coupling leads to shorter timescales.
-On cosmological scales, the primordial quantum perturbations undergo complete decoherence by the end of inflation, thereby transitioning to a classical state.
-In this paper, we require that decoherence occurs sufficiently at the scale of interest, which imposes a lower bound constraint on the interaction parameter.
+The dynamics of an open quantum system are described by the Lindblad master equation, which governs the time evolution of a quantum system interacting with its environment.
+Such environmental couplings induce decoherence, causing the off-diagonal elements of the system’s density matrix to diminish and driving the density matrix towards a diagonal form [41, 42, 43].
+This decoherence modifies primordial perturbations, leading to corrections in their power spectrum.
+The decoherence timescale—over which the off-diagonal elements become negligible—depends on the interaction strength between the system and its environment: stronger coupling results in faster decoherence.
+In this paper, we impose the condition that primordial quantum perturbations at relevant scales undergo complete decoherence by the end of inflation, ensuring that these perturbations become classical thereafter.
+We analyze the impact of decoherence through numerical computation of the modified power spectrum for a polynomial attractor model featuring an ultra-slow-roll stage.
+Furthermore, we derive constraints on the interaction parameter by employing the observational data P-ACT-LB-BK18 from the combination of Planck data, Atacama Cosmology Telescope (ACT) data, CMB lensing, baryon acoustic oscillation (BAO) distance measurements from Dark Energy Spectroscopic Instrument (DESI) and B-mode measurements from the BICEP/Keck telescopes (BK18) [44, 45].
 This paper is organized as follows.
 In Sec.
-II, we present the Lindblad equation that the quantum perturbations satisfy during inflation, along with its exact solution under the assumption of linear interactions.
-In Sec.
-III, we provide the parameters of our model and present the results of numerical calculations regarding the impact of decoherence, as well as constraints on the interaction strength.
-We summarize this paper in Sec.
+II, we present the Lindblad equation along with its exact solution under the assumption of linear interaction.
+In Sec III, we present the results of numerical calculations regarding the impact of decoherence, as well as constraints on the interaction strength.
+The paper is concluded in Sec.
 IV.
@@
 $\hat{H}_{\bm{k}}=\frac{1}{2}\int_{\mathbb{R}^{3}}\mathrm{d}^{3}\bm{k}\left[\omega^{2}(\eta,\bm{k})\hat{v}_{\bm{k}}\hat{v}^{\dagger}_{\bm{k}}+\hat{p}_{\bm{k}}\hat{p}^{\dagger}_{\bm{k}}\right],$ (1)
-where $\hat{v}_{k}$ is the Mukhanov-sasaki variable in Fourier space and $\hat{p}_{k}$ is its conjugate momentum, which obeys $\hat{p}_{k}=\hat{v}_{k}^{\prime}$.
-A prime denotes a derivative with respect to conformal time $\eta$.
+where $\hat{v}_{k}$ is the Mukhanov-sasaki variable in Fourier space, $\hat{p}_{k}=\hat{v}_{k}^{\prime}$ is its conjugate momentum, and a prime denotes a derivative with respect to conformal time $\eta$.
 The frequency of mode $\bm{k}$ is given by
 $\omega^{2}(\eta,\bm{k})=\bm{k}^{2}-\frac{(\sqrt{\epsilon_{1}}a)^{\prime\prime}}{\sqrt{\epsilon_{1}}a},$ (2)
-where $\epsilon_{1}$ is the first slow-roll parameter, $\epsilon_{1}=-\dot{H}/H^{2}$, with $H$ representing the Hubble parameter and a dot indicating differentiation with respect to time.
-The Hamiltonian (1) describes a isolated system.
-However, the primordial perturbations could interact with other degrees of freedom in the early universe, making it more realistic to consider the perturbations as an open quantum system [32, 35].
-For an open quantum system that interacts with the environment, we can describe the total system in the Hilbert space $\mathcal{E}=\mathcal{E}_{S}\otimes\mathcal{E}_{E}$, where the total Hamiltonian is
+[…]
+For an open quantum system that interacts with its environment, the total Hamiltonian is
 $\hat{H}=\hat{H}_{0}+\hat{H}_{\mathrm{int}}=\hat{H}_{S}\otimes\hat{I}_{E}+\hat{I}_{S}\otimes\hat{H}_{E}+g\hat{H}_{\mathrm{int}},$ (3)
-where $H_{S}$ is the Hamiltonian of the system, which acts in $\mathcal{E}_{S}$, and $H_{E}$ is the Hamiltonian of the environment, which acts in $\mathcal{E}_{E}$, $H_{\mathrm{int}}$ is the interaction Hamiltonian and $g$ is the coupling constant.
-Suppose that the system and the environment only couple through local interactions, then the interaction Hamiltonian can be written as
+where $\hat{H}_{S}$ is the Hamiltonian of the quantum system which acts in the Hilbert space $\mathcal{E}_{S}$, $\hat{H}_{E}$ is the Hamiltonian of the environment which acts in the Hilbert space $\mathcal{E}_{E}$, $\hat{H}_{\mathrm{int}}$ is the interaction Hamiltonian and $g$ is the coupling constant.
+Suppose that the quantum system and its environment couple through local interactions only, then the interaction Hamiltonian can be written as
 $\hat{H}_{\mathrm{int}}=\int\mathrm{d}^{3}\bm{x}\hat{A}\otimes\hat{B},$ (4)
 where $\hat{A}$ acts as a local operator within the system’s subsystem, while $\hat{B}$ represents a local operator associated with the environmental component.
-The symbol $\bm{x}$ represents the conformal spatial coordinate.
-According to the reasons mentioned in [32, 35], we can consider $\hat{A}$ to have the form
+Following Refs.
+[31, 34], we consider $\hat{A}$ to have the form
 $\hat{A}=\hat{v}^{n},$ (5)
 where $n$ is a constant.
-By tracing out the environmental degrees of freedom from the total density matrix $\hat{\rho}_{\text{tot}}$, we can derive the reduced density matrix for the system $\hat{\rho}_{S}$
+By tracing out the environmental degrees of freedom from the total density matrix $\hat{\rho}_{\text{tot}}$, the reduced density matrix for the quantum system $\hat{\rho}_{S}$ becomes
 $\hat{\rho}_{S}=\mathrm{Tr}_{E}\,\hat{\rho}_{\text{tot}}.$ (6)
-Using the Born and Markov approximations, we can derive the equation of motion of $\hat{\rho}_{S}$, i.e. the Lindblad equation [41, 42, 32]
+Using the Born and Markov approximations, the Lindblad equation for the reduced density matrix $\hat{\rho}_{S}$ is [46, 41, 31]
 $\frac{\mathrm{d}\hat{\rho}_{S}}{\mathrm{d}\eta}=i\left[\hat{\rho}_{S},\hat{H}_{S}\right]-\frac{\gamma}{2}\int\mathrm{d}^{3}\bm{x}\,\mathrm{d}^{3}{\bm{y}}\,C_{B}\left({\bm{x}},{\bm{y}}\right)\left[\left[\hat{\rho}_{S},\hat{A}(\bm{y})\right],\hat{A}(\bm{x})\right],$ (7)
-where $\gamma=2g^{2}\eta_{c}$ with $\eta_{c}$ being the conformal auto-correlation time of $\hat{B}$, and $C_{B}$ is the same-time correlation function of $\hat{B}$, given by $C_{B}(\bm{x},\bm{y})=\text{Tr}_{E}[\rho_{E}\hat{B}(\eta,\bm{x})\hat{B}(\eta,\bm{y})]$.
-We assume that the time-dependent parameter $\gamma$ depends on the scale factor as a power law [32]
+where $\gamma=2g^{2}\eta_{c}$ with $\eta_{c}$ being the conformal auto-correlation time of $\hat{B}$, and $C_{B}(\bm{x},\bm{y})=\text{Tr}_{E}[\rho_{E}\hat{B}(\eta,\bm{x})\hat{B}(\eta,\bm{y})]$ is the same-time correlation function of $\hat{B}$.
+The time-dependent parameter $\gamma$ is assumed to vary with the scale factor following a power-law form [31],
 $\gamma=\gamma_{*}\left(\frac{a}{a_{*}}\right)^{p},$ (8)
-where * denotes a reference time when the pivot scale $k_{*}=0.05\mathrm{Mpc}^{-1}$ crosses the Hubble radius, and $p$ is a free parameter.
-Assuming the environment to be isotropic and statistically homogeneous, $C_{B}(\bm{x},\bm{y})$ can be defined as a top hat function [32]
+where * denotes a reference time when the pivot scale $k_{*}=0.05\text{ Mpc}^{-1}$ crosses the horizon, and $p$ is a free parameter.
+Assuming the environment to be isotropic and statistically homogeneous, $C_{B}(\bm{x},\bm{y})$ can be written as a top hat function [31]
 $C_{B}(\bm{x},\bm{y})=\bar{C}_{B}\Theta\left(\frac{a|\bm{x}-\bm{y}|}{\ell_{\mathrm{E}}}\right),$ (9)
@@
 $\Theta(x)=\begin{cases}1,&\mathrm{if}\ x<1\\ 0,&\mathrm{otherwise}\end{cases}.$ (10)
-Expanding Eq. (9) in Fourier space, we can obtain
+Expanding Eq. (9) in Fourier space, we obtain
 $\tilde{C}_{B}(\bm{k})\approx\sqrt{\frac{2}{\pi}}\frac{\bar{C}_{B}\ell^{3}_{\mathrm{E}}}{3a^{3}}\Theta\left(\frac{k\ell_{\mathrm{E}}}{a}\right).$ (11)
-Considering the case where the environment consists of a heavy scalar field $\varphi$ with mass $M\gg H$, and the interaction Hamiltonian is
+Considering the case where the environment consists of a heavy scalar field $\varphi$ with mass $M\gg H$, the interaction Hamiltonian is [31]
 $\hat{H}_{\mathrm{int}}=\lambda\mu^{4-n-m}\int\mathrm{d}^{3}\bm{x}\sqrt{-g}\,\hat{\phi}^{n}(\eta,\bm{x})\hat{\varphi}^{m}(\eta,\bm{x}),$ (12)
-where the parameter $\mu$ is defined as a constant mass scale and $\hat{\phi}=\hat{v}/a$. In this case, the parameters can be represented as [32]
+where the parameter $\mu$ is a constant mass scale and $\hat{\phi}=\hat{v}/a$. In this case, we have [31]
 $\bar{C}_{B}$ =$=\left\{\left(2m-1\right)!!-\sigma\left(m\right)\left[\left(m-1\right)!!\right]^{2}\right\}\left(\frac{37}{504\pi^{2}}\frac{H^{6}}{M^{4}}\right)^{m}\,,$ (13)
@@
 where !!double-factorial!!!! denotes the double factorial and
-$\sigma(m)=\begin{cases}1,\quad m\,\,\mathrm{is\,\,even}\\ 0,\quad m\,\,\mathrm{is\,\,odd}\end{cases}.$ (16)
-In order to solve the Lindblad equation (7), we decompose $\hat{v}_{\bm{k}}$ and $\hat{p}_{\bm{k}}$ into their real and imaginary parts
+$\sigma(m)=\begin{cases}1,\quad\text{if }m\,\,\text{is even,}\\ 0,\quad\text{if }m\,\,\text{is odd.}\end{cases}$ (16)
+For the convenience of solving the Lindblad equation (7), we decompose the operators $\hat{v}_{\bm{k}}$ and $\hat{p}_{\bm{k}}$ into their real and imaginary parts,
 $\hat{v}_{\bm{k}}=\frac{1}{\sqrt{2}}\left(\hat{v}_{\bm{k}}^{\mathrm{R}}+i\hat{v}_{\bm{k}}^{\mathrm{I}}\right),\quad\hat{p}_{\bm{k}}=\frac{1}{\sqrt{2}}\left(\hat{p}_{\bm{k}}^{\mathrm{R}}+i\hat{p}_{\bm{k}}^{\mathrm{I}}\right).$ (17)
 For a real operator $\hat{v}(\eta,\bm{x})$, we have $\hat{v}_{-\bm{k}}=\hat{v}^{\dagger}_{\bm{k}}$, which implies $\hat{v}^{\mathrm{R}}_{-\bm{k}}=\hat{v}^{\mathrm{R}}_{\bm{k}}$ and $\hat{v}^{\mathrm{I}}_{-\bm{k}}=-\hat{v}^{\mathrm{I}}_{\bm{k}}$.
-These relations show that not all $\hat{v}_{\bm{k}}$ operators are independent degrees of freedom, we can calculate the variables $v_{\bm{k}}^{\mathrm{R}}$ and $v_{\bm{k}}^{\mathrm{I}}$ in $\bm{k}\in\mathbb{R}^{3+}$.
+These relations show that not all $\hat{v}_{\bm{k}}$ are independent degrees of freedom, we need to calculate the variables $v_{\bm{k}}^{\mathrm{R}}$ and $v_{\bm{k}}^{\mathrm{I}}$ only for $\bm{k}\in\mathbb{R}^{3+}$.
 Under the condition of linear interaction, the free Hamiltonian (1) can be written as
-$\hat{H}_{\bm{k}}^{s}=\frac{1}{2}\int_{\mathbb{R}^{3+}}\mathrm{d}^{3}\bm{k}\left[\omega^{2}(\eta,\bm{k})(\hat{v}_{\bm{k}}^{s})^{2}+(\hat{p}_{\bm{k}}^{s})^{2}\right],$ (18)
+$\hat{H}_{\bm{k}}^{s}=\frac{1}{2}\int_{\mathbb{R}^{3}}\mathrm{d}^{3}\bm{k}\left[\omega^{2}(\eta,\bm{k})(\hat{v}_{\bm{k}}^{s})^{2}+(\hat{p}_{\bm{k}}^{s})^{2}\right],$ (18)
 […]
 $\hat{\rho}_{S}=\prod_{\bm{k}\in\mathbb{R}^{3+}}\prod_{s=\mathrm{R},\mathrm{I}}\hat{\rho}_{\bm{k}}^{s}.$ (19)
-then we can obtain the Lindblad equation in the Fourier space for linear interaction [32]
+The Lindblad equation (7) in the Fourier space for linear interaction becomes [31]
 $\frac{\mathrm{d}\hat{\rho}_{\bm{k}}^{s}}{\mathrm{d}\eta}=-i\left[\hat{\mathcal{H}}_{\bm{k}}^{s},\hat{\rho}_{\bm{k}}^{s}\right]-\frac{\gamma}{2}(2\pi)^{3/2}\tilde{C}_{B}({\bm{k}})\left[\hat{v}_{\bm{k}}^{s},\left[\hat{v}_{\bm{k}}^{s},\hat{\rho}_{\bm{k}}^{s}\right]\right].$ (20)
-Note that Eq. (20) is valid only for linear interaction. We can define a combined interaction parameter as follows
+Note that Eq. (20) is valid only for linear interaction. For the convenience of discussion, we introduce a combined interaction parameter,
 $k_{\gamma}\equiv\sqrt{\frac{8\pi}{3}\bar{C}_{B}{}_{*}\ell_{E}^{3}\frac{\gamma_{*}}{a_{*}^{3}}},$ (21)
 which combines all the parameters of interaction and possesses the dimension of comoving wavenumber.
-Then we define the eigenvectors $|v_{\rm k}^{s}\rangle$ of the operator $\hat{v}_{\rm k}^{s}$, which satisfy equation $\hat{v}_{\rm k}^{s}|v_{\mathbf{k}}^{s}\rangle=v_{\rm k}^{s}|v_{\rm k}^{s}\rangle$.
-By applying the bra $\langle v_{\bm{k}}^{s,(1)}|$ and the ket $|v_{\bm{k}}^{s,(2)}\rangle$ to the Lindblad equation Eq.
-(20), the Lindblad equation can be solved exactly, the solution is given by [32]
-$\left\langle v_{\bm{k}}^{s,(1)}\right|\hat{\rho}_{\bm{k}}^{s}\left|v_{\bm{k}}^{s,(2)}\right\rangle=\frac{\left(2\pi\right)^{-1/2}}{\sqrt{\left|v_{\bm{k}}\right|^{2}+\mathcal{J}_{\bm{k}}}}\mathrm{exp}\left\{-\frac{{v^{s,(2)}_{\bm{k}}}^{2}+{v^{s,(1)}_{\bm{k}}}^{2}+i{\left|v_{\bm{k}}\right|^{2}}^{\prime}\left[{v^{s,(2)}_{\bm{k}}}^{2}-{v^{s,(1)}_{\bm{k}}}^{2}\right]}{4\left(\left|v_{\bm{k}}\right|^{2}+\mathcal{J}_{\bm{k}}\right)}\right\}$
-×$\qquad\qquad\times\exp\Biggl{\{}-\frac{\left[v^{s,(2)}_{\bm{k}}-v^{s,(1)}_{\bm{k}}\right]^{2}}{2\left(\left|v_{\bm{k}}\right|^{2}+\mathcal{J}_{\bm{k}}\right)}\left(\mathcal{I}_{\bm{k}}\mathcal{J}_{\bm{k}}-\mathcal{K}_{\bm{k}}^{2}+\left|v_{\bm{k}}^{\prime}\right|^{2}\mathcal{J}_{\bm{k}}+\left|v_{\bm{k}}\right|^{2}\mathcal{I}_{\bm{k}}-{\left|v_{\bm{k}}\right|^{2}}^{\prime}\mathcal{K}_{\bm{k}}\right)$
-$\qquad\qquad-\frac{i\mathcal{K}_{\bm{k}}}{2\left(\left|v_{\bm{k}}\right|^{2}+\mathcal{J}_{\bm{k}}\right)}\left[{v^{s,(2)}_{\bm{k}}}^{2}-{v^{s,(1)}_{\bm{k}}}^{2}\right]\Biggr{\}},$ (22)
-where $\mathcal{I}_{\bm{k}}$, $\mathcal{J}_{\bm{k}}$ and $\mathcal{K}_{\bm{k}}$ are defined by
+To solve the Lindblad equation, we define the eigenvectors $|v_{\rm k}^{s}\rangle$ of the operator $\hat{v}_{\rm k}^{s}$ satisfying equation $\hat{v}_{\rm k}^{s}|v_{\mathbf{k}}^{s}\rangle=v_{\rm k}^{s}|v_{\rm k}^{s}\rangle$.
+Projecting the Lindblad equation Eq.
+(20) with the bra $\langle v_{\bm{k}}^{s_{1}}|$ and the ket $|v_{\bm{k}}^{s_{2}}\rangle$, we get the the solution [31]
+$\left\langle v_{\bm{k}}^{s_{1}}\right|\hat{\rho}_{\bm{k}}^{s}\left|v_{\bm{k}}^{s_{2}}\right\rangle=\frac{\left(2\pi\right)^{-1/2}}{\sqrt{\left|v_{\bm{k}}\right|^{2}+\mathcal{J}_{\bm{k}}}}\mathrm{exp}\left\{-\frac{\left({v^{s_{2}}_{\bm{k}}}\right)^{2}+\left({v^{s_{1}}_{\bm{k}}}\right)^{2}+i{\left|v_{\bm{k}}\right|^{2}}^{\prime}\left[\left({v^{s_{2}}_{\bm{k}}}\right)^{2}-\left({v^{s_{1}}_{\bm{k}}}\right)^{2}\right]}{4\left(\left|v_{\bm{k}}\right|^{2}+\mathcal{J}_{\bm{k}}\right)}\right\}$
+×$\qquad\qquad\times\exp\Biggl{\{}-\frac{\left[v^{s_{2}}_{\bm{k}}-v^{s_{1}}_{\bm{k}}\right]^{2}}{2\left(\left|v_{\bm{k}}\right|^{2}+\mathcal{J}_{\bm{k}}\right)}\left(\mathcal{I}_{\bm{k}}\mathcal{J}_{\bm{k}}-\mathcal{K}_{\bm{k}}^{2}+\left|v_{\bm{k}}^{\prime}\right|^{2}\mathcal{J}_{\bm{k}}+\left|v_{\bm{k}}\right|^{2}\mathcal{I}_{\bm{k}}-{\left|v_{\bm{k}}\right|^{2}}^{\prime}\mathcal{K}_{\bm{k}}\right)$
+$\qquad\qquad-\frac{i\mathcal{K}_{\bm{k}}}{2\left(\left|v_{\bm{k}}\right|^{2}+\mathcal{J}_{\bm{k}}\right)}\left[\left({v^{s_{2}}_{\bm{k}}}\right)^{2}-\left({v^{s_{1}}_{\bm{k}}}\right)^{2}\right]\Biggr{\}},$ (22)
+where $\mathcal{I}_{\bm{k}}$, $\mathcal{J}_{\bm{k}}$ and $\mathcal{K}_{\bm{k}}$ are defined as
 $\mathcal{I}_{\bm{k}}\left(\eta\right)$ $\equiv 4\left(2\pi\right)^{3/2}\int_{-\infty}^{\eta}\mathrm{d}\eta^{\prime}\gamma\left(\eta^{\prime}\right)\tilde{C}_{B}\left(\bm{k},\eta^{\prime}\right)\mathrm{Im}^{2}\left[v_{\bm{k}}\left(\eta^{\prime}\right){v_{\bm{k}}^{*}}^{\prime}\left(\eta\right)\right]\,,$ (23)
@@
 $\mathcal{K}_{\bm{k}}\left(\eta\right)$ $\equiv 4\left(2\pi\right)^{3/2}\int_{-\infty}^{\eta}\mathrm{d}\eta^{\prime}\gamma\left(\eta^{\prime}\right)\tilde{C}_{B}\left(\bm{k},\eta^{\prime}\right)\mathrm{Im}\left[v_{\bm{k}}\left(\eta^{\prime}\right){v_{\bm{k}}^{*}}^{\prime}\left(\eta\right)\right]\mathrm{Im}\left[v_{\bm{k}}\left(\eta^{\prime}\right){v_{\bm{k}}^{*}}\left(\eta\right)\right]\,,$ (25)
-where $v_{\bm{k}}(\eta)$ is the solution of the Mukhanov-Sasaki equation.
+and $v_{\bm{k}}(\eta)$ is the solution of the Mukhanov-Sasaki equation.
 Using Eq.
@@
 where we take $M_{\mathrm{Pl}}^{2}=1$.
-On the other hand, for linear interaction, the tensor power spectrum remains unchanged by the corrections in the Lindblad equation [35].
+For linear interaction, the tensor power spectrum remains unchanged by the corrections introduced in the Lindblad equation [34].
 Therefore, we can utilize the spectral index and the tensor-to-scalar ratio to constrain $k_{\gamma}$ for linear interaction.
 The interaction with the environment tends to suppress the off-diagonal elements of the density matrix when expressed in the basis of the eigenstates of the interaction operator [41, 31].
-From the exact solution of the density matrix Eq.
-(II), we can study decoherence during inflation.
-The decoherence parameter $\delta_{\bm{k}}$ defined in [32] represents the reduction in the off-diagonal elements of the density matrix due to environmental interactions, and is defined as follows
+From the exact solution to Eq.
+(II) satisfied by the density matrix, we can study decoherence during inflation.
+The parameter $\delta_{\bm{k}}$ defined in [31] represents the reduction in the off-diagonal elements of the density matrix due to environmental interactions, and is defined as follows
 $\delta_{\bm{k}}(\eta)\equiv\mathcal{I}_{\bm{k}}\mathcal{J}_{\bm{k}}-\mathcal{K}_{\bm{k}}^{2}+\left|v_{\bm{k}}^{\prime}\right|^{2}\mathcal{J}_{\bm{k}}+\left|v_{\bm{k}}\right|^{2}\mathcal{I}_{\bm{k}}-{\left|v_{\bm{k}}\right|^{2}}^{\prime}\mathcal{K}_{\bm{k}}\,.$ (28)
 It is obvious that $\delta_{\bm{k}}$ is related to $k_{\gamma}$.
-When $\delta_{\bm{k}}\gg 1$, decoherence is expected to be complete.
-We determine that by the end of inflation, the decoherence of the scales of interest should be finished, which in turn constrains $k_{\gamma}$.
-In the study referenced as [32], the researchers took into account the constant $\epsilon_{1}$ and the constant Hubble parameter $H$.
-In this paper, we will compute the results for a concrete model with an ultra-slow-roll stage.
-We compare the results of our model’s numerical computations with the analytical results reported in [32].
-Then, we constraint $k_{\gamma}$ using the polynomial attractor model.
+When $\delta_{\bm{k}}\gg 1$, decoherence is expected to be fully established.
+The decoherence process associated with cosmological scales of interest must have been completed by the end of the inflation, thereby establishing constraint on the interaction parameter $k_{\gamma}$.
+In [31], analytical expressions for the corrections to the power spectrum were derived within the slow-roll approximation.
+In this paper, we numerically calculate the correction to the power spectrum for a polynomial attractor model featuring an ultra-slow-roll stage, and compare the result with the analytical result reported in [31].
 III Models and numerical results
-We choose the modified polynomial attractor model in [45]
+We choose the modified polynomial attractor model with the potential [47]
 $V(\psi)=V_{0}[1+c_{4}f^{-1}(\psi)+c_{1}f(\psi)+c_{2}f^{2}(\psi)+c_{3}f^{3}(\psi)]^{2},$ (29)
-where $f(\psi)=1/\psi^{3}$, the parameters of the model are given by Table. 1.
-From Eq. (8) and (15), we can see that for linear interaction,
+where $f(\psi)=1/\psi^{3}$, and the parameters of the model are given in Table. 1.
+From Eqs. (8) and (15), we see that for linear interaction,
 $\gamma(\eta^{\prime})=\gamma_{*}\left(\frac{a(\eta^{\prime})}{a_{*}}\right)^{5}.$ (30)
-From Eq. (11) and (13), we have
+From Eqs. (11) and (13), we have
 $\tilde{C}_{B}(\bm{k},\eta^{\prime})=\left(\frac{H(\eta^{\prime})}{H_{*}}\right)^{6m}\left(\frac{a_{*}^{3}}{a(\eta^{\prime})^{3}}\right)\sqrt{\frac{2}{\pi}}\frac{\bar{C}_{B}^{*}\ell_{E}^{3}}{3a_{*}^{3}}\Theta\left(\frac{k\ell_{E}}{a(\eta^{\prime})}\right),$ (31)
@@
 $\bar{C}_{B}^{*}=\left\{\left(2m-1\right)!!-\sigma\left(m\right)\left[\left(m-1\right)!!\right]^{2}\right\}\left(\frac{37}{504\pi^{2}}\frac{H_{*}^{6}}{M^{4}}\right)^{m}\,.$ (32)
-Combine Eq. (24), (30)and (31), we have
+Combining Eqs. (24), (30) and (31), we have
 $4(2\pi)^{3/2}\gamma(\eta^{\prime})\tilde{C}_{B}(\bm{k},\eta^{\prime})$ $=2\left(\frac{a(\eta^{\prime})}{a_{*}}\right)^{2}\left(\frac{H(\eta^{\prime})}{H_{*}}\right)^{6m}\frac{8\pi}{3}\frac{\bar{C}_{B}^{*}\ell_{E}^{3}\gamma_{*}}{a_{*}^{3}}\Theta\left(\frac{k\ell_{E}}{a(\eta^{\prime})}\right)$
 $=2\left(\frac{a(\eta^{\prime})}{a_{*}}\right)^{2}\left(\frac{H(\eta^{\prime})}{H_{*}}\right)^{6m}k_{\gamma}^{2}\,\Theta\left(\frac{k\ell_{E}}{a(\eta^{\prime})}\right).$ (33)
-The remaining sector of Eq. (24) can be written as
+and
 $\mathrm{Im}^{2}\left[v_{\bm{k}}(\eta^{\prime})v_{\bm{k}}^{*}(\eta)\right]=$$\left\{\mathrm{Im}\left[\left(v^{\mathrm{R}}_{\bm{k}}(\eta^{\prime})+iv^{\mathrm{I}}_{\bm{k}}(\eta^{\prime})\right)\left(v^{\mathrm{R}}_{\bm{k}}(\eta)-iv^{\mathrm{I}}_{\bm{k}}(\eta)\right)\right]\right\}^{2}$
 $=$ $\left(v^{\mathrm{R}}_{\bm{k}}(\eta^{\prime})\right)^{2}\left(v^{\mathrm{I}}_{\bm{k}}(\eta)\right)^{2}+\left(v^{\mathrm{I}}_{\bm{k}}(\eta^{\prime})\right)^{2}\left(v^{\mathrm{R}}_{\bm{k}}(\eta)\right)^{2}$
-$-2v^{\mathrm{R}}_{\bm{k}}(\eta^{\prime})v^{\mathrm{I}}_{\bm{k}}(\eta^{\prime})v^{\mathrm{R}}_{\bm{k}}(\eta)v^{\mathrm{I}}_{\bm{k}}(\eta),$ (34)
-We take $m=3$ so that $\mu$ in Eq.
-(15) vanishes.
-In order to compare with [32], we take $k_{\gamma}=5\times 10^{-5}\text{Mpc}^{-1}$ and $H_{*}\ell_{E}=10^{-3}$.
+$-2v^{\mathrm{R}}_{\bm{k}}(\eta^{\prime})v^{\mathrm{I}}_{\bm{k}}(\eta^{\prime})v^{\mathrm{R}}_{\bm{k}}(\eta)v^{\mathrm{I}}_{\bm{k}}(\eta).$ (34)
+To simplify the calculation, we take $m=3$ so that $\gamma$ defined in Eq.
+(15) is independent of $\mu$.
+In order to make comparison with the results in [31], we take $k_{\gamma}=5\times 10^{-5}\text{ Mpc}^{-1}$ and $H_{*}\ell_{E}=10^{-3}$.
 The top hat function in Eq.
-(III) can be written as
+(III) is
 $\Theta\left(\frac{k\ell_{E}}{a(\eta^{\prime})}\right)=\Theta\left(\frac{k}{a(\eta^{\prime})H(\eta^{\prime})}\frac{H(\eta^{\prime})}{H_{*}}H_{*}\ell_{E}\right).$ (35)
-Then we can calculate the modified power spectrum numerically.
-Models $V_{0}$ $\delta$ $c_{1}$ $c_{4}$ $\psi_{*}$ $N$ $n_{s}$ $r$ $k_{\text{peak}}$
-A1 8.33×$\times 10^{-10}$ 1.1007×$\times 10^{-3}$1.07 3.2×$\times 10^{-4}$ 3.75 63.2 0.9658 0.025 $3.0\times 10^{8}$
-Table 1: Parameters of the model A1.
-Figure 1: The black dashed line is the modified power spectrum, the red dotted line is the slow-roll result of $\Delta P_{k}$, the blue solid line is the numerical result of $\Delta P_{k}$ calculated from model A1.
-We calculate the modified term $\Delta P_{k}$ of the attractor model A1.
-The results of the modified power spectrum and $\Delta P_{k}$ are shown in Fig.
-1.
+With these chosen parameters, we calculate the modified power spectrum numerically and the results are shown in Fig. 1.
+Model $V_{0}$ $\delta$ $c_{1}$ $c_{4}$ $\psi_{*}$ $N$ $n_{s}$ $r$ $k_{\text{peak}}$
+A 5.3×$\times 10^{-10}$ 9.81×$\times 10^{-4}$1.1 2.3×$\times 10^{-4}$ 4 51.4 0.9743 0.017 $8.7\times 10^{10}$
+Table 1: Parameters of the model A.
+Figure 1: The power spectrum for the polynomial attractor model A.
+The black solid line is the modified power spectrum, the red dotted line is the correction to the power spectrum $\Delta P_{k}$ within slow-roll approximation, and the black dashed line is the numerical result for the correction to the power spectrum $\Delta P_{k}$.
 From Fig.
-1, we see that at the pivot scale $k_{*}$, both the slow-roll model and ultra-slow-roll model have the same value of $\Delta P_{k}$.
+1, we see that at the pivot scale $k_{*}$, the numerical result of $\Delta P_{k}$ for the ultra-slow-roll model agrees with analytical one within the slow-roll approximation.
+This is because the pivot scale crosses the horizon during the slow-roll stage.
 However, as the Hubble parameter gradually decreases, $\Delta P_{k}$ also decreases.
-We also find that at the minimum of the power spectrum, $\Delta P_{k}$ has a peak in the order of $10^{-6}$, but the correction near the peak is negligible ($\sim 10^{-8}$).
-This means that the correction due to decoherence almost does not change the abundance of primordial black holes and the scalar induced gravitational waves.
-Figure 2: $n_{s}$ and $r$ with respect to $k_{\gamma}$. The red solid lines is the numerical results from model A1, the black dashed lines is the analytical results for slow-roll.
+At the minimum of the power spectrum, $\Delta P_{k}$ exhibits a peak on the order of $10^{-4}$; yet, the correction near this peak remains negligible (approximately $\sim 10^{-8}$).
+This indicates that the decoherence-induced correction has an almost negligible effect on the abundance of primordial black holes and the scalar-induced secondary gravitational waves.
+Figure 2: The behavior of the observables $n_{s}$ and $r$ with respect to $k_{\gamma}$.
+The red solid lines are the numerical results and the black dashed lines are the analytical results within slow-roll approximation.
 Figure 3: The spectral index and tensor-to-scalar ratio vary with $k_{\gamma}$.
-The red solid line is the numerical result, the black dashed line is the analytical result, the black dash dot line is the restriction of $n_{s}$ and $r$ at $1\sigma$ and $2\sigma$ from Planck18.
-As $k_{\gamma}$ increases, the spectral index and tensor-to-scalar ratio decrease.
-Figure 4: Decoherence parameter $\delta_{\bm{k}}$ as a function of e-folds $N$ at $k=k_{*}$.
-The black dashed line is the result of slow-roll model, the red dotted line is the result of model A1.
-The vertical solid line indicates the e-folds value at the starting point of $\delta_{k}$ evolution.
-From Fig.
-1, we find that the correction at pivot scale is more significant, which means that we can restrict $k_{\gamma}$ using the spectral tilt $n_{s}$ and the tensor-to-scalar ratio $r$ at pivot scale.
-The analytical results for $n_{s}$ and $r$ in the case of linear interaction are [32]
+The red solid line represents the numerical results, the black dashed line corresponds to the analytical results within slow-roll approximation, and the black solid lines indicate the $1\sigma$ and $2\sigma$ constraints from P-ACT-LB-BK18.
+As shown by the lines, $n_{s}$ and $r$ decrease as $k_{\gamma}$ increases.
+Figure 4: Decoherence parameter $\delta_{\bm{k}}$ at $k=k_{*}$ as a function of $N-N_{*}$.
+The red dotted line represents the numerical results, and the black dashed line corresponds to the analytical results within slow-roll approximation.
+The vertical solid line marks the e-fold value at which the evolution of $\delta_{k}$ begins.
+Now we restrict $k_{\gamma}$ using observational constraints on the spectral tilt $n_{s}$ and the tensor-to-scalar ratio $r$ at the pivot scale.
+The analytical results for $n_{s}$ and $r$ in the case of linear interaction within slow-roll approximation are [31]
 $n_{s}=n_{s}|_{\mathrm{standard}}-\frac{\alpha}{1+\alpha}(6m-2)\epsilon_{1*},$ (36)
 and
-$r=\frac{r|_{\mathrm{standard}}}{1+\alpha},\quad\alpha=\frac{\pi}{6}\frac{k_{\gamma}^{2}}{k_{*}^{2}}.$ (37)
-We perform numerical computations to determine the relationships between $k_{\gamma}$, $n_{s}$ and $r$.
-The results are shown in Fig.
+$r=\frac{r|_{\mathrm{standard}}}{1+\alpha},$ (37)
+where $\alpha=\pi k_{\gamma}^{2}/(6k_{*}^{2})$.
+By varying $k_{\gamma}$, we numerically calculate $n_{s}$ and $r$ at the pivotal scale and the results are shown in Fig.
 2.
-We can see that when $k_{\gamma}>10^{-2}\,\text{Mpc}^{-1}$, the results begin to change significantly, when $k_{\gamma}>0.2\text{Mpc}^{-1}$, the changes in the results are minimal.
-The tensor-to-scalar ratios are well-aligned, whereas the spectral tilts deviate at large $k_{\gamma}$.
-We plot the results in the $n_{s}-r$ plane to restrict the upper bound of $k_{\gamma}$, the results are shown in Fig.
-3.
-Starting from the initial point, as $k_{\gamma}$ grows, both $n_{s}$ and $r$ diminish.
-When $k_{\gamma}=0.026\,\text{Mpc}^{-1}$ and $k_{\gamma}=0.042\,\text{Mpc}^{-1}$, the results are within the $1\sigma$ and $2\sigma$ confidence intervals of the observation, respectively.
-This is the upper bound of $k_{\gamma}$ given by model A1.
-We then calculate the decoherence parameter $\delta_{\bm{k}}$ at pivot scale and compare with the result in [32].
-The results are shown in Fig.
+In the range $0.2\,\text{Mpc}^{-1}>k_{\gamma}>10^{-2}\,\text{Mpc}^{-1}$, both $n_{s}$ and $r$ vary significantly; outside this range, they remain nearly constant.
+For the tensor-to-scalar ratio $r$, the numerical result agrees with the analytical one within slow-roll approximation.
+However, for the scalar spectral tilt $n_{s}$, deviations appear when $k_{\gamma}>10^{-2}\,\text{Mpc}^{-1}$.
+In Fig.
+3, we plot the results in $n_{s}-r$ plane along with the $1\sigma$ and $2\sigma$ constraints by P-ACT-LB-BK18.
+From Fig.
+3, it is evident that both $n_{s}$ and $r$ decrease as $k_{\gamma}$ increases.
+The $1\sigma$ and $2\sigma$ constraints from P-ACT-LB-BK18 give upper bounds $k_{\gamma}<0.041\,\text{Mpc}^{-1}$ and $k_{\gamma}<0.061\,\text{Mpc}^{-1}$, respectively.
+To constrain $k_{\gamma}$ using decoherence, we present the results for the decoherence parameter $\delta_{\bm{k}}$ at the pivot scale in Fig.
 4.
-Since model A1 is in the slow-roll stage at the pivot scale, the analytical and numerical results are almost the same.
-After $N-N_{*}=4$, $\delta_{\bm{k}}$ increase exponentially.
-We require that for the scales $k_{*}$ and $k_{\mathrm{peak}}$, decoherence is complete.
-For model A1, this imposes a constraint on $k_{\gamma}$ such that $k_{\gamma}$ must be much greater than $10^{-28}\,\text{Mpc}^{-1}$.
-This is a rather relaxed constraint.
+Since the pivotal scale crosses the horizon during the slow-roll phase, the numerical results agree well with the analytical approximations.
+When $N-N_{*}\geq 4$, where $N$ is the number of e-folds before the end of inflation and $N_{*}$ is the number of e-folds remaining at the horizon crossing of the pivot scale, $\delta_{\bm{k}}$ increases exponentially.
+We require decoherence to be complete for both the scales $k_{*}$ and $k_{\mathrm{peak}}$ by the end of inflation, leading to the lower bound $k_{\gamma}\gg 10^{-17}\,\text{Mpc}^{-1}$.
+Compared to the results obtained in Ref.
+[31], our bound is tighter because we demand that perturbations not only at CMB scales but also at the peak of the power spectrum fully decohere by the end of inflation.
 IV Conclusion
-In this paper, we examine the effects of decoherence on the power spectrum of a specific inflationary model that incorporates an ultra-slow-roll phase.
-Decoherence arises from the interaction of a quantum system with its environment and is described by the Lindblad equation.
-Based on the solutions to the Lindblad equation for linear interactions provided by J.
-Martin et al.
-[32], we numerically computed the corrections to the power spectrum of a polynomial attractor model due to quantum decoherence and constrained the interaction parameters $k_{\gamma}$ using observations.
-We found that decoherence significantly alters the power spectrum only on large scales, and it is consistent with previous work at the pivot scales.
-As the scale decreases, the corrections to the power spectrum due to decoherence also diminish, and at the trough of the power spectrum, the corrections will result in a peak.
-We also calculated the relationship between the spectral index and the tensor-to-scalar ratio at the CMB with $k_{\gamma}$, and using the observational data from Planck 2018, we determined an upper limit for $k_{\gamma}$ of $0.042$  Mpc-1.
-We calculate the decoherence parameter $\delta_{k}$, which is used to characterize the degree of completion of decoherence; decoherence is considered to be complete when $\delta_{k}\gg 1$.
-We computed the evolution of $\delta_{k}$ over time at the pivot scale, and the results are roughly consistent with the analysis of the slow-roll regime.
-We require that decoherence be completed at the peak scale of the power spectrum by the end of inflation, which will provide a lower limit for the interaction parameter $k_{\gamma}$.
-Our model suggests a lower limit of $10^{-28}\,\text{Mpc}^{-1}$.
-Since exact solutions are only provided for linear interactions, the aforementioned numerical calculations must satisfy the conditions of linear interaction.
+In this paper, we examine the effects of decoherence on the power spectrum in an inflationary model featuring an ultra-slow-roll phase.
+Decoherence, arising from interactions between the quantum system and its environment, is described by the Lindblad equation.
+Previous studies, employing the slow-roll approximation, calculated the interaction’s impact on the power spectrum and placed preliminary constraints on both the interaction parameter $k_{\gamma}$ and decoherence parameter $\delta_{\bm{k}}$.
+Using solutions to the Lindblad equation for linear interactions from Ref.
+[31], we numerically compute the corrections to the power spectrum of a polynomial attractor model and constrain the interaction parameter $k_{\gamma}$ using observational data.
+We find that decoherence significantly modifies the power spectrum primarily on large scales, consistent with analytical results obtained under the slow-roll approximation.
+As the scale decreases, the corrections become smaller, but a peak appears at the trough of the power spectrum.
+By analyzing the impact of $k_{\gamma}$ on observables such as the scalar spectral tilt $n_{s}$ and the tensor-to-scalar ratio $r$, we establish an upper limit of $k_{\gamma}<0.061$  Mpc-1 from the P-ACT-LB-BK18 data.
+We also calculate the decoherence parameter $\delta_{k}$ at the pivot scale, which is used to characterize the degree of completion of decoherence.
+Requiring complete decoherence at both the pivot scale $k_{*}$ and the peak scale $k_{\mathrm{peak}}$ by the end of inflation leads to a tighter lower bound, $k_{\gamma}\gg 10^{-17}\,\text{Mpc}^{-1}$.
+In conclusion, we constrain the interaction parameter within the range $10^{-17}\text{Mpc}^{-1}<k_{\gamma}<0.061\text{Mpc}^{-1}$.
+Acknowledgements.
+This work is supported in part by the National Key Research and Development Program of China under Grant No. 2020YFC2201504.